OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12?

OpenStudy (uri):

Formula y=mx+b So arrange the equation first 6y=-4x+12

OpenStudy (uri):

y=-4/6x+12/6 y=-2/3x+2 slope of line is -2/3x

OpenStudy (anonymous):

Which coordinate grid shows the graph of y = 2x – 3?

OpenStudy (anonymous):

I wouldn't suggest including x in the answer though. Slope is =m and m=-2/3, not -2/3x

OpenStudy (anonymous):

@iviliciousz, close this question and post your next question in new thread

OpenStudy (anonymous):

What is the y-intercept for the graph of 3x – 5y = 6?

OpenStudy (anonymous):

**see previous post

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i need help What is the y-intercept for the graph of 3x – 5y = 6?

OpenStudy (anonymous):

What is the slope of the line 8x + 6y = 12?